IONS AS PROBES OF SUB-PS WATER NETWORK DYNAMICS.

\underline{G. W. SCHWAAB}, D. A. SCHMIDT, S. FUNKNER, B. BORN, M. HAVENITH ^{a}, Department of Physical Chemistry II, Ruhr-University Bochum, D-44780 Bochum, Germany; O. BIRER, Department of Chemistry, Koç University, 34450 Istanbul, Turkey; R. GNANASEKARAN, and D. M. LEIT-NER, Department of Chemistry, University of Nevada, 89557 Reno, Nevada, USA.

We present THz measurements of salt solutions which shed new light on the controversy of salts as kosmotropes (structure makers) or chaotropes (structure breakers). We performed concentration dependent narrow-band THz-absorption spectroscopy on fifteen alkali-halide salt solutions around 85 cm^{−1} (2.5 THz) and wide-band (30-300 cm^{−1}) THz Fourier transform measurements on six alkali halide salt solutions. All solutions show an increased THz-absorption compared to pure water with a linear concentration dependence. Our comprehensive data set is well-described by a model including damped harmonic oscillations (rattling modes) of both anions and cations within the water network. This model well-predicts key features of THz spectra for a variety of salt solutions. Complementary molecular dynamics simulations using the TIP3P water model support experiments and show that the fast sub-ps ionic motions and their surroundings are almost decoupled. These findings provide a complete description of the solute-induced changes in the THz solvation dynamics for the investigated salts and suggests a treatment of the ions as simple defects in an H-bond network. Our results show that THz spectroscopy is a powerful experimental tool to establish a new insights on contributions to the structuring of water by anions and cations.

aThe authors acknowledge financial support by the VW Stiftung Az I/84 302, the BMBF (grant 05 KS7PC2), FOR 618 and NSF (grant CHE-0910669).